OpenStudy (anonymous):

Choose the equation of the line passing through the point (-2, 6) and parallel to y = one halfx - 8. A. y = one halfx + 7 B. y = one halfx + 5 C. y = 2x + 7 D. y = 2x + 5

OpenStudy (anonymous):

Hint: parallel lines have same slope.

OpenStudy (anonymous):

what about perpindicular lines?

OpenStudy (anonymous):

the product of the slopes of perpindicular lines equal to -1 if two lines are perpendicular with slopes m1 and m2 then \[\Large m_{1}m_{2}=-1\]

OpenStudy (anonymous):

Parallel lines have same slope. Perpendicular lines have will have the reverse slope of each other. I forgot what it was called lol.

OpenStudy (anonymous):

For example, Parallel would be something like y = 2x + 5 and y = 2x - 9, but Perpendicular would be something like y = 2x + 5 and y = - 1/2 x + 5